IRRATIONALITY MEASURES OF log 2 AND π/\sqrt{3}

NICOLAS BRISEBARRE

Abstract. Using a class of polynomials that generalizes Legendre polynomials, we unify previous works of E. A. Rukhadze, A. K. Dubitskas, M. Hata, D. V. and G. V. Chudnovsky about irrationality measures of log 2 and π/\sqrt{3}.